You break a piece of Styrofoam packing material, and it releases lots of little spheres whose electric charge makes them stick annoyingly to you.

If two of the spheres carry equal charges and repel with a force of 21mN when they're 14mm apart, what's the magnitude of the charge on each?


please make it clear so i would understand.


thanks
Answer #1

I have solved this question earlier with different figures. Please workout using yours figures. If you need any further help just PM me. If I have helped you please rate me 5 stars first (before you rate anyone else).

You break a piece of Styrofoam packing material, and it releases lots of little spheres whose electric charge makes them stick annoyingly to you.

If two of the spheres carry equal charges and repel with a force of 22 { mN} when they're 16 { mm} apart, what's the magnitude of the charge on each?


we can consider spheres as point charges

hence electrostatic force between them= kq2/r2

hence 22*10-3 = 9*109 q2/.0162

hence charge on the spheres = 0.25 nC   
Answer #2

we can consider spheres as point charges

hence electrostatic force between them= kq2/r2

hence 22*10-3 = 9*109 q2/.0162

hence charge on the spheres = 0.25 nC   
Answer #3

F=kQq/r^2
k=1/(4*pi*8.85*10^-12).
And q=Q:

q=sqrt(F*r^2/k)

q=sqrt(21*10^-3*(0.014)^2*4*pi*8.85*10^-12)

q= 2.19*10^-8 C
To two sigfig:
q=2.2*10^-8 C
Answer #4

F= kq1*q2/r^2 = k*q^2/r^2 (since q1=q2=q)

=> 21*10^-3 = 8.99*10^9*q^2/(14*10^-3)

=> q = sqrt((21*10^-3)*(14*10^-3)/(8.99*10^9)) = 1.80*10^-7 C

)